OpenStudy (anonymous):

The length of a rectangle is 6 1/2 inches and the width is 3 3/4 inches. What is the ratio, using whole numbers, of the length to the width? A. 26:30 B. 13:15 C. 15:26 D. 26:15 Basically, how do you do proportions? ;)

OpenStudy (mertsj):

\[\frac{6.5}{3.75}=\frac{650}{375}\]

OpenStudy (mertsj):

Reduce the fraction.

OpenStudy (anonymous):

okay.. so do you know how to convert mixed fractions into improper fractions?

OpenStudy (anonymous):

the point is for you to understand proportions.. not to just give you answers! :)

OpenStudy (anonymous):

6 1/2 = 13/2.... and 3 3/4 = 15/4.

OpenStudy (anonymous):

Yes cupcake1123. 13/2 & 15/4

OpenStudy (anonymous):

so when they say length to width.. length always goes on the top of the fraction, and width in the denominator. so divide length by width.. (13/2)/(15/4)

OpenStudy (anonymous):